OpenStudy (avanti):

Trig Question - Solve: sin^2 (theta) + cos (theta) =2 (Hint: Use the Pythagorean identity, sin^2 (theta) +cos^2 (theta) = 1 , to replace sin^2(theta) in the given equation). I still don't quite know what to do or how to show work, please help!

OpenStudy (psymon):

Right. So if sin^2(x) + cos^2(x) = 1, you could also say either sin^2(x) is the same as 1 - cos^2(x) or that cos^2(x) is the same as 1-sin^2(x). That's what theyre hinting at you to do.

OpenStudy (avanti):

so are they asking me to find theta?

OpenStudy (psymon):

Nah. Just that you can do this: 

OpenStudy (avanti):

so would sin^2 =1 - cos^2(x) be my answer?

OpenStudy (anonymous):

I think they do want you to find theta.

OpenStudy (psymon):

Well eventually, yes, but just that he can turn sin^2(x) into 1-cos^2(x), allowing him to do another quadratic solving thing to eventually get theta.

OpenStudy (psymon):

Sorry if I made it confusing.

OpenStudy (anonymous):

I don't think so... \[\left( 1-\cos^2\theta \right)+\cos \theta=2\] Is where you are no go from there.

OpenStudy (avanti):

I'm not sure how to find theta from that

OpenStudy (anonymous):

Well I would substitute a variable say u in place of cos then simplify. So what would you have if you subbed u in for cos?

OpenStudy (avanti):

(1-u^2 θ) + u θ =2 ? then factor?

OpenStudy (anonymous):

Well close but remember theta is the argument of cos so you really arent replacing cos but cos theta. so (1-u^2) + u = 2

OpenStudy (anonymous):

Are you sure that the original equation was \[\sin^2 \theta + \cos \theta = 2\]?

OpenStudy (psymon):

I dont think it was, because there is no solution if that is the original.

OpenStudy (avanti):

This is what it looks like

OpenStudy (avanti):

look in the top left corner

OpenStudy (anonymous):

Yep because the greatest value of cos theta will be 1 and that only occurs when theta is ... 0, 2pi, 4pi,... and the max value of sin^2 theta is 1 as well and that occurs when theta is pi/2, 3pi/2, 5pi/2 ... so these two values will never add to two so the final answer is no solution. If it is as written.

OpenStudy (anonymous):

\[\sin^2 (\theta) + \cos (\theta) =2\] \[1-\cos^2(\theta )+\cos(\theta)=2\] \[-\cos^2(\theta)+\cos(\theta)=1\] \[\cos^2(\theta)-\cos(\theta)+1=0\]

OpenStudy (psymon):

Unless satellite has some magic, I say no solution, too.

OpenStudy (anonymous):

solve \[x^2-x+1=0\] using the quadratic formula

OpenStudy (anonymous):

then take the inverse cosine of the result

OpenStudy (anonymous):

That would yield a complex number so in there is no solution in the real numbers.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

lol so it will!

OpenStudy (psymon):

Haha, damn, no magic.

OpenStudy (anonymous):

yeah no solutions

OpenStudy (anonymous):

in fact you know there are no solutions from the start

OpenStudy (anonymous):

No solutions in the Real numbers

OpenStudy (anonymous):

you have \[\sin^2(\theta)+\cos(\theta)=2\]

OpenStudy (anonymous):

That is what I said! see above

OpenStudy (anonymous):

the largest \(\cos(\theta)\) can be is 1

OpenStudy (anonymous):

Yes as i said above.

OpenStudy (anonymous):

and also the largest \(\sin^2(\theta)\) can be is 1 and they are not one at the same place

OpenStudy (anonymous):

yes as I said above.

OpenStudy (psymon):

Probably the reason they gave you the pythagorean hint, just so you could see that you're going to find a way to get 2 out of that.

OpenStudy (anonymous):

yeah so i see i am a little slow

OpenStudy (psymon):

*not going to

OpenStudy (anonymous):

Thanks for the support. :-)

OpenStudy (avanti):

yeah i was wondering why they gave me the Pythagorean identity

OpenStudy (avanti):

well thank you all so so much for helping me! :)

OpenStudy (anonymous):

I think it was a poorly designed problem.

OpenStudy (psymon):

A very implicit way to prove a point I guess :/

OpenStudy (avanti):

yeah i totally agree

OpenStudy (psymon):

Back to the math cave.